Grade 12th PassPhysical Chemistry

0.3g of an organic compound containing C,H,O on combustion yields 0.44g of CO2 and 0.18g of H2O,with two oxygen atoms per molecule
.FIND EMPRICAL FORMULA AND MOLECULAR FORMULA OF COMPOUND

AVANISH KUMAR GUPTA
8 Years agoGrade 12th Pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
 
First write the eqn and balance it.
CxHyOz           +        (2x + y/2) O2            --------->>>>>           xCO2           +        y/2 H2O
moles of CO2 = 0.44/(12x + 32x)
moles of H2O = 0.18/9y
moles of compound = 0.3/(12x + y + 16z)
Now U Can easily use mole mole analysis and solve the eqns for x y and z.
 
 
Regards
Arun (askIITians forum expert)